OpenStudy (anonymous):

Convert the radical below into a fractional exponent The square root of 4^3

OpenStudy (anonymous):

4 ^ 3/2

OpenStudy (anonymous):

\[\sqrt{4^{3}}\] imagine that it actually looks like this ...\[\sqrt[2]{4^{3}}\] get rid of the \[\sqrt{}\] and the 2 goes under the 3 so you get 4^ 3/2

OpenStudy (anonymous):

Yup yup

OpenStudy (anonymous):

i forget the actually name of the "2" but even if it looked like this \[\sqrt[5]{}\] the 5 would go under the 3

OpenStudy (anonymous):

\[\sqrt[5]{8}\] i dentify the radicand in the radical below

OpenStudy (anonymous):

radicand is the number or function under the radical, so 8 in this case.

OpenStudy (anonymous):

When dealing with fractional exponents there is a little trick to make things go faster. The numerator of a fractional exponent tells that you will be taking some number to the power of the numerator. And the denominator tells you what root you will take. Here is an example \[x^\frac{1}{2}\] This can be written as \[\sqrt{x^1}\] Which is just the square root of x. So in your problem \[\sqrt{4^3}\] You are cubing the 4. So the three is the numerator of the fractional exponent. You are taking the square root so a 2 will be in the denominator. So your exponent will be \[\frac{3}{2}\] Leaving the 4 alone you will get the answer of \[4^\frac{3}{2}\]